# PrefixQuant: Static Quantization Beats Dynamic through Prefixed Outliers in LLMs

### v1

@@ 5.2 Block-wise Fine-tuning
Recent studies demonstrate that block-wise fine-tuning (Shao et al., 2023; Chen et al., 2024a) enhances performance by considering inter-layer interactions (Li et al., 2021). We initialize all quantization parameters using grid search (Lin et al., 2023; 2024b) and then fine-tune each transformer block with mean square error loss sequentially. For trainable parameters, we follow EfficientQAT (Chen et al., 2024a) by activating the training of all quantization parameters and original full-precision weights. Additionally, unlike dynamic activation quantization, the offline pre-computed quantization parameters of static activation quantization are inherently trainable. To maintain simplicity, we use block-wise quantization in this work and leave the end-to-end finr-tuning of EfficientQAT (Chen et al., 2024a) for future performance improvements.
| Model | Method | Quant Type | Wiki PPl | Avg. Acc. |
| Llama-2-7B | FP16 | - | 5.47 | 69.04 |
[…]
| QuaRot | dynamic | 6.19 | 64.69 | |
| DuQuant | dynamic | 6.20 | 66.25 | |
| SpinQuant | dynamic | 5.95 | 65.35 | |
| PrefixQuant w/o FT | static | 6.22 | 66.84 | |
| PrefixQuant | static | 6.01 | 66.37 | |
| Llama-2-13B | FP16 | - | 4.88 | 71.73 |
[…]
| QuaRot | dynamic | 5.45 | 69.01 | |
| DuQuant | dynamic | 5.39 | 69.13 | |
| SpinQuant | dynamic | 5.24 | 69.24 | |
| PrefixQuant w/o FT | static | 5.50 | 69.92 | |
| PrefixQuant | static | 5.32 | 70.36 | |
| Llama-2-70B | FP16 | - | 3.32 | 76.72 |
[…]
| QuaRot | dynamic | 3.83 | 75.43 | |
| DuQuant | dynamic | 3.77 | 74.75 | |
| SpinQuant | dynamic | 3.70 | 75.19 | |
| PrefixQuant w/o FT | static | 4.41 | 73.29 | |
| PrefixQuant | static | 3.81 | 75.48 | |
| Llama-3-8B | FP16 | - | 6.14 | 72.71 |
[…]
| QuaRot | dynamic | 8.41 | 65.15 | |
| DuQuant | dynamic | 8.14 | 67.13 | |
| SpinQuant | dynamic | 7.36 | 68.23 | |
| PrefixQuant w/o FT | static | 7.93 | 68.37 | |
| PrefixQuant | static | 7.43 | 71.08 | |
| Llama-3-70B | FP16 | - | 2.85 | 80.03 |
[…]
| DuQuant | dynamic | 5.67 | 74.89 | |
| PrefixQuant w/o FT | static | 5.23 | 76.40 | |
| PrefixQuant | static | 4.41 | 77.18 | |
| * Grayed results use Wikitext2 as calibaration dataset. | ||||
| Atom apply 128 group size quantization to both weight and activations. | ||||
@@ 6.1 Setups
Baseline. PrefixQuant is a versatile method applicable to any precision. We conduct experiments on three precisions: W8A8KV8, W4A8KV4, and W4A4KV4. In PrefixQuant, weight uses per-channel symmetric quantization. KV cache uses per-head symmetric static quantization for 4-bit and per-tensor symmetric static quantization for 8-bit. Activation (inputs of linear layers) uses per-tensor static quantization. We compare PrefixQuant with QuaRot (Ashkboos et al., 2024b), Atom (Zhao et al., 2023), DuQuant (Lin et al., 2024a), QoQ (Lin et al., 2024b), SmoothQuant (Xiao et al., 2023a) and SpinQuant (Liu et al., 2024a). Following QoQ, we reproduce all these methods except SpinQuant with Pile (Gao et al., 2020) calibration dataset to avoid over-fitting for fair comparisons. The detailed quantization configuration and results sources of these comparison methods can be found at Sec. B. Note that all comparison methods use dynamic quantization if without specific mentioned, and would suffer dramatic performance degeneration likes “+ static quantization” in Table 6.
@@
Models and datasets. We evaluate PrefixQuant on the Llama-2, Llama-3, Llama-3-Instruct families, Mistral-7B-v0.3, and Qwen-2-7B models. Following previous literature (Shao et al., 2023; Lin et al., 2024b), we assess PrefixQuant quantized models on language modeling and zero-shot tasks. Specifically, we evaluate on WikiText2 (Merity et al., 2016) with a 2048 context length for perplexity, and on PIQA (Bisk et al., 2020), ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), and WinoGrande (Sakaguchi et al., 2021) using lm_eval v0.4.2 (Gao et al., 2024). For accuracy, we report acc for WinoGrande and acc_norm for HellaSwag, Arc_Challenge, Arc_Easy, and PIQA, following Qserve (Lin et al., 2024b)†††Some weight-only quantization works such as EfficientQAT (Chen et al., 2024a) and QuiP# (Tseng et al., 2024) report acc for all tasks..
@@
Grid Search Setting. For all experiments with static quantization, we initialize the quantization parameters through grid search on 8 Pile (Gao et al., 2020) samples with a 1024 sequence length. We minimize the layer outputs for fine-grained quantization (per-channel/per-head) and block outputs for per-tensor quantization. In the performance comparison tables, “PrefixQunt w/o FT” indicates finishing the quantization only with grid search and without fine-tuning.
@@ 6.2 Comparison Results
Results on W4A4KV4. Table 3 shows the comparison results for W4A4KV4. PrefixQuant with static quantization significantly outperforms the previous state-of-the-art QuaRot, which uses dynamic quantization. For instance, in Llama-3-8B, PrefixQuant without fine-tuning surpasses QuaRot by in WikiText2 perplexity and points in average accuracy. Fine-tuning further improves these results to in WikiText2 perplexity and points in average accuracy.
| Model | Method | Activation Quant | Wiki PPl | Avg. Acc. |
| Llama-2-7B | QoQ | dynamic | 5.75 | 67.22 |
| QuaRot | dynamic | 5.73 | 67.11 | |
| PrefixQuant w/o FT | static | 5.76 | 67.86 | |
| PrefixQuant | static | 5.68 | 68.90 | |
| Llama-2-13B | QoQ | dynamic | 5.12 | 70.56 |
| QuaRot | dynamic | 5.07 | 69.96 | |
| PrefixQuant w/o FT | static | 5.08 | 71.07 | |
| PrefixQuant | static | 5.07 | 71.25 | |
| Llama-2-70B | QoQ | dynamic | 3.52 | 75.91 |
| QuaRot | dynamic | 3.46 | 76.31 | |
| PrefixQuant w/o FT | static | 3.60 | 75.00 | |
| PrefixQuant | static | 3.50 | 76.50 | |
| Llama-3-8B | QoQ | dynamic | 6.89 | 71.35 |
| QuaRot | dynamic | 6.80 | 71.68 | |
| PrefixQuant w/o FT | static | 6.90 | 70.29 | |
| PrefixQuant | static | 6.62 | 72.46 | |
| Llama-3-70B | QoQ | dynamic | 4.36 | 78.12 |
| QuaRot | dynamic | 3.73 | 78.92 | |
| PrefixQuant w/o FT | static | 3.55 | 77.82 | |
| PrefixQuant | static | 3.43 | 78.70 |
Results on W4A8KV8. Table 4 presents the comparison results for W4A8KV8. Without fine-tuning, PrefixQuant performs comparably to QoQ (Lin et al., 2024b). After fine-tuning, PrefixQuant outperforms both QoQ and QuaRot in most models. For instance, PrefixQuant surpasses QoQ (Lin et al., 2024b) by perplexity and accuracy points in Llama-3-8B.
Results on W8A8KV8. Table 18 includes the comparison with various methods in W8A8KV8 quantization. We can find that SmoothQuant, QuaRot, and PrefixQuant all attain near lossless performance without fine-tuning. Notably, our PrefixQuant is unique in employing static quantization, which enhances inference efficiency. Additionally, earlier methods like CushionCache (Son et al., 2024) and QFeP (Yang et al., 2024), despite also using prefixed tokens in the KV cache to support coarser quantization, exhibit marked performance decline even under W8A8 as illustrated in Table 17.
@@
Results on weight-only quantization. PrefixQuant can also improve existing weight-only quantization methods by reducing outlier noise in MSE loss calculations. As shown in Table 16, PrefixQuant enhances the average accuracy by and points on W2A16g128 Llama-3-8B and Llama-3-70B, respectively, based on the state-of-the-art uniform quantization method EfficientQAT (Chen et al., 2024a). See Sec. G for more details.
| Batchsize | 1 | 4 |
| on a RTX 3090 GPU (ms) | ||
@@ 6.3 Inference Speed
In this section, we evaluate the end-to-end inference speed of PrefixQuant in the W4A4 quantization scenario. We do not consider KV quantization here because it saves memory footprint through more computation overhead and only achieves speedup with large batch sizes (Liu et al., 2024b). Table 5 shows the speedup of W4A4 quantization in the prefilling stage. Our PrefixQuant improves the QuaRot speedup from to on the A100-80GB GPU, and from to on the RTX 3090 GPU. In Sec. D, we also provide comprehensive apple-to-apple comparisons of sub-modules, such as quantization kernels and quantized linears, demonstrating the significant superiority of PrefixQuant over the existing dynamic quantization approach QuaRot (Ashkboos et al., 2024b).
| Method | Activation Quant. | W8A8KV8 | W4A8KV4 | W4A4KV4 |
| QuaRot | dynamic | 6.17 | 6.75 | 8.33 |
| RTN | dynamic | 6.26 | 12.66 | 1282.34 |
| + rotation | dynamic | 6.17 | 10.88 | 24.98 |
| + grid search | dynamic | 6.17 | 8.91 | 16.47 |
| + static quantization | static | 7.27 | 29.07 | 141.02 |
| + prefixed outliers | static | 6.17 | 6.90 | 7.93 |
| + block-wise fine-tuning | static | 6.17 | 6.63 | 7.41 |
6.4 Ablation Studies
We examine the impact of various quantization techniques implemented in PrefixQuant. Our analysis starts with W4A4KV4 round-to-nearest (RTN) quantization on Llama-3-8B, involving per-channel weight quantization, per-token dynamic activation quantization, and per-head dynamic KV-cache quantization. We apply different techniques step-by-step and report the WikiText2 perplexity in Table 6. We find that both Hadamard rotation and grid search initialization improve performance. Then, perplexity collapses due to static quantization of activation and KV cache, but introducing prefixed outliers significantly recovers performance, even surpassing results before introducing static quantization. These benefits arise not only by reducing information loss from outlier tokens but also by helping to find accurate quantization parameters in grid searches through isolating extremely large outliers () in activation. Additionally, block-wise fine-tuning improves performance except on W8A8KV8, which is nearly lossless without fine-tuning. More ablation results related to the training dataset, training epochs, dynamic quantization, the number of prefixed tokens, and the content of prefixed tokens are in Sec. F in the Appendix.
@@ 7 Conclusion
We propose PrefixQuant, which enables static quantization to outperform dynamic quantization by effectively handling token-wise outliers through a novel prefixing approach. This technique also stabilizes model training, making it a plug-and-play module that enhances the performance of other optimization-based methods. The simplicity and broad applicability of PrefixQuant make it a promising direction for future LLM compression and optimization research.
Appendix A Reproducibility Statement
In this section, we summarize the necessary information to reproduce our results. First, PrefixQuant is based on Hadamard rotation, as detailed in Sec.C. Our main contribution, setting prefixed outliers, is discussed in Sec.5.1. After configuring prefixed outliers in the KV-cache, we initialize the quantization parameters using grid search. We also offer optional block-wise fine-tuning to enhance performance. Detailed setups for grid search and fine-tuning are available in Sec.6.1. Additionally, we provide the source of detailed results for each compared method in Sec.B.
Appendix B Configuration and Data Sources of Comparison Methods
Quantization Configurations. In this study, we establish the quantization granularity for each comparison method based on the specifications provided in the original papers. Details on these settings are given in Table 7.
| Method | Weight | Activation | KV Cache |
| SmoothQuant | per-channel symmetric | per-token symmetric dynamic | per-tensor symmetric static |
| Atom | per-group symmetric | per-group symmetric dynamic | per-group asymmetric dynamic |
| QoQ | per-channel asymmetric | per-token symmetric dynamic | per-group asymmetric dynamic |
| QuaRot | per-channel symmetric | per-token dynamic symmetric | per-group asymmetric dynamic |
| DuQuant | per-channel asymmetric | per-token dynamic asymmetric | per-token asymmetric dynamic |
| PrefixQuant | per-channel symmetric | per-tensor symmetric static | per-head symmetric static |
Data Sources. We compare our proposed PrefixQuant with several other methods: QuaRot (Ashkboos et al., 2024b), Atom (Zhao et al., 2023), QoQ (Lin et al., 2024b), SmoothQuant (Xiao et al., 2023a), SpinQuant (Liu et al., 2024a), and EfficientQAT (Chen et al., 2024a). The data for our comparisons either come directly from the official publications of these methods, from other papers, or from our own reproduction of the methods. The source of the data for each method is outlined as follows:
@@ Appendix B Configuration and Data Sources of Comparison Methods
Appendix C Details of Rotation
Hadamard rotation (Ashkboos et al., 2024b; Liu et al., 2024a) redistributes outlier channels across all channels, achieving uniform distribution within each token. The Hadamard matrix is an orthogonal matrix with , and its entries are at the same scale. Hadamard rotation can be applied to all activations and use inverse rotation on corresponding weights to maintain computational invariance (Ashkboos et al., 2024a). Specifically, the rotation includes absorbable and online rotations. As shown in Figure 6, we follow SpinQuant (Liu et al., 2024a) to set , , and rotations, details as follows.

### v2

@@ 4.3 Block-wise Fine-tuning
Recent studies show that block-wise fine-tuning (Shao et al., 2023; Chen et al., 2024a) improves performance by accounting for inter-layer interactions (Li et al., 2021). To further enhance the performance of quantized models, we fine-tune each transformer block sequentially using a mean squared error (MSE) loss. Specifically, we introduce trainable parameters for the activation quantizer to balance the rounding and clipping errors in quantization. For dynamic activation quantization in PrefixQuant-O1, we set the tensor-wise clipping factors as trainable. Note that the clipping factors cannot be token-wise, as long-context scenarios introduce excessive storage overhead with token-wise clipping factors. For static activation quantization in PrefixQuant-O2, the quantization parameters (scaling factors and zero-points) are inherently trainable. For weight quantization, we follow the approach of EfficientQAT (Chen et al., 2024a), enabling the training of all weights and weight quantization parameters.
@@ 5.1 Setups
Baseline. PrefixQuant is a versatile method applicable to any precision. We conduct experiments on two mainstream precisions: W4A8KV4, and W4A4KV4. The detailed quantization settings are illustrated in Table 1. PrefixQuant-O1 is consistent with existing methods for fair comparisons, and PrefixQuant-O2 targets to push the limitation of more efficient static quantization. We compare PrefixQuant with QuaRot (Ashkboos et al., 2024b), Atom (Zhao et al., 2023), DuQuant (Lin et al., 2024a), QoQ (Lin et al., 2024b), and SpinQuant (Liu et al., 2024b). Following QoQ, we reproduce all these methods except SpinQuant with Pile (Gao et al., 2020) calibration dataset to avoid over-fitting for fair comparisons. The detailed quantization configuration and results sources of these comparison methods can be found at Sec. A.
| Model | Method | PPL | Acc. |
| 2-7B | FP16 | 5.47 | 69.04 |
[…]
| QuaRot | 6.19 | 64.69 | |
| DuQuant | 6.20 | 66.25 | |
| SpinQuant | 5.95 | 65.35 | |
| PrefixQuant-O1 | 5.93 | 66.74 | |
| PrefixQuant-O2 | 6.01 | 66.37 | |
| 2-13B | FP16 | 4.88 | 71.73 |
[…]
| QuaRot | 5.45 | 69.01 | |
| DuQuant | 5.39 | 69.13 | |
| SpinQuant | 5.24 | 69.24 | |
| PrefixQuant-O1 | 5.24 | 70.05 | |
| PrefixQuant-O2 | 5.32 | 70.36 | |
| 2-70B | FP16 | 3.32 | 76.72 |
[…]
| QuaRot | 3.83 | 75.43 | |
| DuQuant | 3.77 | 74.75 | |
| SpinQuant | 3.70 | 75.19 | |
| PrefixQuant-O1 | 3.62 | 76.23 | |
| PrefixQuant-O2 | 3.81 | 75.48 | |
| 3-8B | FP16 | 6.14 | 72.71 |
[…]
| QuaRot | 8.41 | 65.15 | |
| DuQuant | 8.14 | 67.13 | |
| SpinQuant | 7.36 | 68.23 | |
| PrefixQuant-O1 | 7.26 | 71.31 | |
| PrefixQuant-O2 | 7.43 | 71.08 | |
| 3-70B | FP16 | 2.85 | 80.03 |
[…]
| DuQuant | 5.67 | 74.89 | |
| PrefixQuant-O1 | 4.16 | 77.08 | |
| PrefixQuant-O2 | 4.41 | 77.18 |
| Model | Method | PPL | Acc. |
| 2-7B | FP16 | 5.47 | 69.04 |
[…]
| QuaRot | 5.73 | 67.11 | |
| PrefixQuant-O1 | 5.67 | 68.04 | |
| PrefixQuant-O2 | 5.68 | 68.09 | |
| 2-13B | FP16 | 4.88 | 71.73 |
[…]
| QuaRot | 5.07 | 69.96 | |
| PrefixQuant-O1 | 5.05 | 71.25 | |
| PrefixQuant-O2 | 5.07 | 71.25 | |
| 2-70B | FP16 | 3.32 | 76.72 |
[…]
| QuaRot | 3.46 | 76.31 | |
| PrefixQuant-O1 | 3.44 | 76.82 | |
| PrefixQuant-O2 | 3.50 | 76.50 | |
| 3-8B | FP16 | 6.14 | 72.71 |
[…]
| QuaRot | 6.80 | 71.68 | |
| PrefixQuant-O1 | 6.59 | 72.57 | |
| PrefixQuant-O2 | 6.62 | 72.46 | |
| 3-70B | FP16 | 2.85 | 80.03 |
[…]
| QuaRot | 3.73 | 78.52 | |
| PrefixQuant-O1 | 3.37 | 78.50 | |
| PrefixQuant-O2 | 3.43 | 78.70 |
| Method | Precision | MMLU Acc. |
| - | FP16 | 62.07 |
| QuaRot | W4A4KV4 | 34.25 |
| DuQuant | W4A4KV4 | 50.77 |
| SpinQuant | W4A4KV4 | 51.93 |
| PrefixQuant-O1 | W4A4KV4 | 56.00 |
| PrefixQuant-O2 | W4A4KV4 | 54.65 |
| QuaRot | W4A8KV4 | 38.37 |
| DuQuant | W4A8KV4 | 58.01 |
| SpinQuant | W4A8KV4 | 58.25 |
| PrefixQuant-O1 | W4A8KV4 | 60.49 |
| PrefixQuant-O2 | W4A8KV4 | 59.20 |
@@
Models and datasets. We evaluate PrefixQuant on the Llama-2, Llama-3, Llama-3-Instruct families, Mistral-7B-v0.3, and Qwen-2-7B models. Following previous literature (Shao et al., 2023; Lin et al., 2024b), we assess PrefixQuant quantized models on language modeling and zero-shot tasks. Specifically, we evaluate on WikiText2 (Merity et al., 2016) with a 2048 context length for perplexity, and on 5 zero-shot reasoning tasks, including PIQA (Bisk et al., 2020), ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), and WinoGrande (Sakaguchi et al., 2021). We also test models on more challenge zero-shot MMLU (Hendrycks et al., 2020). All accuracy are measured through lm_eval v0.4.2 (Gao et al., 2024). For accuracy, we report acc for WinoGrande and acc_norm for HellaSwag, Arc_Challenge, Arc_Easy, and PIQA, following Qserve (Lin et al., 2024b).
| Method | Activation Quant. | W4A8KV4 | W4A4KV4 | |
| QuaRot | dynamic | 6.75 | 8.33 | |
| Base | RTN | dynamic | 12.66 | 1282.34 |
| + rotation | dynamic | 10.88 | 24.98 | |
| + Grid search initialization | dynamic | 9.08 | 11.70 | |
[…]
| + block-wise fine-tuning | dynamic | 6.59 | 7.23 | |
[…]
| + prefixed outliers | static | 6.90 | 7.93 | |
| + block-wise fine-tuning | static | 6.62 | 7.41 |
@@
Grid Search Initialization Setting. We initialize the quantization parameters through grid search on 8 Pile (Gao et al., 2020) samples with a 1024 sequence length. We minimize the layer outputs for fine-grained quantization (per-channel/per-head) and block outputs for per-tensor quantization.
@@ 5.2 Comparison Results
Results on W4A4KV4. Table 2 presents the comparison results for W4A4KV4. PrefixQuant consistently outperforms existing methods. For example, under the same dynamic quantization setting on Llama-3-8B, PrefixQuant-O1 achieves a WikiText perplexity improvement and points accuracy over DuQuant. Additionally, the more efficient PrefixQuant-O2 for static quantization also surpasses DuQuant, with a perplexity improvement and points accuracy.
Results on W4A8KV8. Table 3 shows the comparison results for W4A8KV8. PrefixQuant-O1 and PrefixQuant-O2 outperform both QoQ and QuaRot across most models. For instance, PrefixQuant-O1 surpasses QoQ (Lin et al., 2024b) by perplexity and points accuracy on Llama-3-8B. Similarly, PrefixQuant-O2 maintains performance benefits with a perplexity improvement and points accuracy.
@@
Results on MMLU. Table 4 presents the comparison results on zero-shot MMLU using Llama-3-8B. PrefixQuant-O1 and PrefixQuant-O2 outperform SpinQuant by and accuracy, respectively, in W4A8KV4 quantization. The performance advantage is even more pronounced in W4A4KV4 quantization, with improvements of and accuracy, respectively.
5.3 Ablation Studies
We analyze the effects of various quantization techniques implemented in PrefixQuant. These techniques are applied incrementally, and the WikiText2 perplexity results are presented in Table 5. The analysis begins with round-to-nearest (RTN) quantization on Llama-3-8B, incorporating Hadamard rotation and grid search initialization. Using W4A4KV4 as an example, we observe that introducing prefixed outliers significantly improves performance. Specifically, perplexity decreases from to for PrefixQuant-O1 and from to for PrefixQuant-O2. These improvements result not only from mitigating information loss caused by outlier tokens but also from enabling more accurate quantization parameter selection during grid searches initialization by isolating extremely large outliers (e.g., values exceeding ) in activations. Additionally, block-wise fine-tuning further enhances performance, reducing perplexity by for PrefixQuant-O1 and by for PrefixQuant-O2 in W4A4KV4 quantization. Additional ablation results, including analyses of the training dataset, training epochs, dynamic quantization, the number of prefixed tokens, and the content of prefixed tokens, are provided in Sec. D of the Appendix.
| Method | prefilling | decoding |
| (ms) | (token/s) | |
| FP16 | 489 | 43 |
@@ 5.4 Inference Speed
In this section, we evaluate the end-to-end inference speed of PrefixQuant under the W4A4 quantization scenario. KV quantization is not considered because it reduces memory usage at the cost of increased computation overhead and only provides speedup with large batch sizes (Liu et al., 2024a). As shown in Table 6, PrefixQuant achieves an approximate speedup in prefilling and a speedup in decoding compared to the FP16 model.
@@ 6 Conclusion
In this paper, we propose PrefixQuant, which provides a comprehensive exploration of outlier tokens and introduces an efficient and effective method to handle them by prefixing these tokens in the KV cache. Additionally, we design new trainable parameters for activation quantization to minimize quantization error. The proposed PrefixQuant method achieves excellent performance across various models, quantization precisions, and granularities. The simplicity and broad applicability of PrefixQuant make it a promising direction for future research on LLM compression and optimization.
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Results Sources of Comparison Methods
We compare our proposed PrefixQuant with several other methods: QuaRot (Ashkboos et al., 2024b), Atom (Zhao et al., 2023), QoQ (Lin et al., 2024b), SmoothQuant (Xiao et al., 2023a), SpinQuant (Liu et al., 2024b), and EfficientQAT (Chen et al., 2024a). The data for our comparisons either come directly from the official publications of these methods, from other papers, or from our own reproduction of the methods. The source of the data for each method is outlined as follows:
@@ Appendix A Results Sources of Comparison Methods
Appendix B Details of Rotation
Hadamard rotation (Ashkboos et al., 2024b; Liu et al., 2024b) redistributes outlier channels across all channels, achieving uniform distribution within each token. The Hadamard matrix is an orthogonal matrix with , and its entries are at the same scale. Hadamard rotation can be applied to all activations and use inverse rotation on corresponding weights to maintain computational invariance (Ashkboos et al., 2024a). Specifically, the rotation includes absorbable and online rotations. As shown in Figure 7, we follow SpinQuant (Liu et al., 2024b) to set , , and rotations, details as follows.